# Arab Spring’s Impact on Science through the Lens of Scholarly Attention, Funding, and Migration

### v3

@@ Discussion
We evaluate the accuracy of the identification process through three tasks: first, assessing whether our method can detect country mentions; second, determining how accurate and comprehensive the method is in identifying these mentions; and third, evaluating the accuracy of detecting mentions of the focal country in this study. To achieve this, we created three samples of articles, totaling 1,000 articles, with titles and abstracts obtained for analysis. The first sample comprises 520 randomly selected papers from our analyzed time frame, with 20 papers chosen from each of the 27 subject areas, including those with and without country mentions. The second sample comprises 280 papers that mention countries without restrictions on subject areas. The third sample comprises 200 randomly selected papers that mention at least one of the 10 target countries analyzed in this study, with 20 papers mentioning each country.
Each paper was independently annotated by two coders, who are among the authors of this paper. They identified all countries or regions mentioned in the titles or abstracts. Coders were instructed not to annotate instances where country names appeared but did not refer to geographical locations. Examples of such exclusions include country names used as part of terms (e.g., Jordan matrix, Congo red), names referring to non-country locations (e.g., New Mexico, Chad Basin), and instances referring to nationalities rather than countries (e.g., Swedish, Iranian) and languages (Georgian, Vietnamese, French). The inter-coder agreement rate was 89.1%, reflecting a high level of consistency in identifying geographical references. Then, two baselines are implemented: one derived from the intersection of the sets of countries found by the two annotators and the other from their union. We use two key measures commonly used in this domain: Accuracy  (?) and Average Jaccard Similarity.
@@
where $n$ is the total number of instances, $L_{i}^{d}$ represents the data labels for the text $i$, $L_{i}^{c}$ represents the labels predicted by the model, for instance, $i$, and $I$ is the indicator function that equals 1 if $L_{i}^{d}=L_{i}^{c}$ and 0 otherwise. A high value in this metric indicates consistent, perfect predictions for all labels in each instance. This metric is highly stringent, requiring a perfect match between predicted and actual label sets for each instance.
@@
Our identification algorithm is evaluated against three established state-of-the-art models: ChatGPT-4.0 (?), BERT-base NER, and BERT-large NER (?).
@@
As a name entity recognition tool, you are responsible for analyzing academic texts from various sources. Your task is to find all the country names mentioned in the text. Constraint: Answer with only the Python list of countries in ISO3 format mentioned in the text that is most accurate and nothing else. Text:text
As shown in Table S3 in the Supplementary materials, our algorithm outperforms all other models, except for those mentioning target countries, where ChatGPT (?) slightly outperforms it. In this task, our dictionary-based approach offers superior general accuracy and is significantly more efficient in terms of computation speed and cost. Running a large language model, such as ChatGPT, on millions of articles would currently cost tens of thousands of dollars and would take anywhere from days to weeks, depending on the implementation.
@@
Since many publications feature multiple authors with multiple affiliations, we employ fractional counting to ensure that each author’s contribution is accurately represented. For example, consider a dummy paper authored by Jane Smith and John Doe. Suppose Jane Smith is affiliated with Country A and Country C, meaning her contribution is split equally (0.5 for each country). In contrast, John Doe, who is only affiliated with Country C, contributes a full count.
@@
ICSR lab extracts funding information, when present, from the funding or acknowledgment sections of research papers. The funding information includes the country of origin for the founding agencies we use to construct the scholarly funding network, and in a similar manner to the scholarly attention network (see Fig. 9). This process involves two steps: constructing a bipartite network and subsequently projecting it into a unipartite network using fractional counting.
In the bipartite network construction, two distinct sets of nodes are defined. On one side are the countries of the funding agencies, and on the other are those mentioned in the publication’s title or abstract. The bipartite network is converted into a unipartite, directed network using fractional counting. In this structure, countries serve as nodes, and the edges, $F^{i}_{j}$, represent the fractionally adjusted counted number of publications that both mention country $i$ and are funded by agencies from country $j$.
@@
In Figure 2, we conducted a bibliometric analysis to assess the extent to which the surge in scholarly attention received in target countries is attributed to research related to the Arab Spring. Scopus  (?) is a bibliometric database that allows users to filter publications by submitting information through a ‘query.’
@@
A search of the Scopus database, which draws from the same database as ICSR, using the query above, conducted in January 2025, yielded 3,458 articles. The same country identification algorithm was applied to extract the country names mentioned in the titles and abstracts of these articles, and 67.8% of the articles referenced at least one country.
@@
To identify the Arab Spring’s impact on scholarly attention and funding, we adopt a Difference-in-Differences (DiD) statistical approach typically used to mimic an experimental research design when only observational data is available. This approach enables us to estimate the average treatment effect, the Arab Spring, by comparing temporal changes in attention/funding to target countries (treatment group) with the rest of the world or unaffected MENA countries (control groups).
We acknowledge that using the rest of the world or relatively unaffected MENA countries as a control group may not provide an ideal counterfactual. However, we aim to highlight indicative patterns rather than precise causal estimates. We emphasize treatment effects’ direction and statistical significance and avoid comparing effect sizes across models.
@@
where the parameter $\beta^{p}_{s}$ captures the difference in trends between treated and control groups before the treatment period. The parameter $\lambda$ reflects the overall time trend common to all countries, while $\theta$ measures the effect of the time-varying covariates $X_{it}$. The term $\gamma_{i}$ accounts for country-specific fixed effects, capturing characteristics that are constant over time but vary across countries. The test’s null hypothesis ($\beta^{p}_{s}=0$) indicates no difference in pre-treatment trends between treated and control countries for subgroup $s\in S$. The alternative hypothesis ($\beta^{p}_{s}\neq 0$) implies a violation of the parallel trends assumption. We conduct the tests at a significance level of $\alpha=0.05$. When the null hypothesis is rejected ($p<\alpha$), it suggests that treated and control countries exhibited different trends before treatment, which undermines the validity of the DiD framework. When the alternate hypothesis is rejected ($p\geq\alpha$), it does not prove that the assumption of parallel trends holds, but it provides statistical support for it. We find that the null hypothesis is only rejected for one coefficient in Table 1 (indicated with a dagger) and in 8 coefficients in Table S5 in Supplementary Materials. We report all $p$-values for these tests in Table S4 in the Supplementary Materials.
@@ Acknowledgments
Table S4: $p$-values for pre-treatment parallel trend test Models control for log GDP, population, and research population where applicable. Each colored column represents $p$-values from separate regressions comparing the target countries with two control groups: the rest of the world and MENA countries less affected by the Arab Spring.

### v4

@@ Discussion
We evaluate the accuracy of the identification process through three tasks: first, assessing whether our method can detect country mentions; second, determining how accurate and comprehensive the method is in identifying these mentions; and third, evaluating the accuracy of detecting mentions of the focal country in this study. To achieve this, we created three samples of articles, totaling 1,000 articles, with titles and abstracts obtained for analysis. The first sample comprises 520 randomly selected papers from our analyzed time frame, with 20 papers chosen from each of the 27 subject areas, including those with and without country mentions. The second sample comprises 280 papers that mention countries without restrictions on subject areas. The third sample includes 200 randomly selected papers that mention at least one of the 10 target countries analyzed in this study, with 20 documents mentioning each country.
Each paper was independently annotated by two coders, who are among the authors of this paper. They identified all countries or regions mentioned in the titles or abstracts. Coders were instructed not to annotate instances where country names appeared but did not refer to geographical locations, following the algorithm’s criteria. The inter-coder agreement rate was 89.1%, reflecting a high level of consistency in identifying geographical references. Then, two baselines are implemented: one derived from the intersection of the sets of countries found by the two annotators and the other from their union. We use two key measures commonly used in this domain: Accuracy/ Exact Match Ratio (?) and Average Jaccard Similarity.
@@
Where $n$ is the total number of instances, $L_{i}^{d}$ represents the data labels for the text $i$, $L_{i}^{c}$ represents the labels predicted by the model, for instance, $i$, and $I$ is the indicator function that equals 1 if $L_{i}^{d}=L_{i}^{c}$ and 0 otherwise. A high value in this metric indicates consistent, perfect predictions for all labels in each instance. This metric is highly strict, requiring a perfect match between predicted and actual label sets for each instance.
@@
Our identification algorithm is evaluated against three established baselines: specialized transformer models such as BERT-base NER and BERT-large NER (?), as well as the large language model ChatGPT-4.0 (?).
@@
As a named entity recognition tool, you are responsible for analyzing academic texts from various sources. Your task is to find all the country names mentioned in the text. Constraint: Answer with only the Python list of countries in ISO 3 format mentioned in the text that is most accurate, and nothing else. Text:text
As shown in Table S5 in the Supplementary Materials, our algorithm outperforms all other models, except for those mentioning target countries, where ChatGPT (?) slightly outperforms it. In this task, our dictionary-based approach offers superior general accuracy and is significantly more efficient in terms of computation speed and cost. Running a large language model, such as ChatGPT, on millions of articles would currently cost tens of thousands of dollars and take anywhere from days to weeks, depending on the implementation.
@@
Since many publications have multiple authors with multiple affiliations, we employ fractional counting to ensure that each author’s contribution is accurately represented. For example, consider a dummy paper authored by Jane Smith and John Doe. Suppose Jane Smith is affiliated with Country A and Country C, meaning her contribution is split equally (0.5 for each country). In contrast, John Doe, who is only affiliated with Country C, contributes a full count.
@@
ICSR lab extracts funding information, when present, from the funding or acknowledgment sections of research papers. The funding information includes the country of origin for the founding agencies. We use this information to construct the scholarly funding network similarly to the scholarly attention network (see Fig. 9). This process involves two steps: constructing a bipartite network and subsequently projecting it into a unipartite network using fractional counting.
In the bipartite network construction, two distinct sets of nodes are defined. On one side are the countries of the funding agencies, and on the other are those mentioned in the publication’s title or abstract. The bipartite network is converted into a unipartite, directed network using fractional counting. In this structure, countries serve as nodes, and the edges, $F^{i}_{j}$, represent the fractionally counted number of publications that mention country $i$ and are funded by agencies from country $j$.
@@
In Figure 2, we conducted an analysis to assess the extent to which the surge in scholarly attention received in target countries is attributed to research related to the Arab Spring. Scopus (?) is a bibliometric database that allows users to filter publications by submitting information through a ‘query.’
@@
A search of the Scopus database, which draws from the same database as ICSR, using the query above, conducted in January 2025, yielded 3,458 articles. The same country identification algorithm was applied to extract the country names mentioned in the titles and abstracts of these articles, and 67.8% of the articles mentioned at least one country.
@@
To identify the Arab Spring’s impact on scholarly attention and funding, we adopt a Difference-in-Differences (DiD) statistical approach typically used to mimic an experimental research design when only observational data is available. This approach enables us to estimate the average treatment effect of the Arab Spring by comparing temporal changes in attention/funding to target countries (treatment group) with those observed in appropriate control groups, namely, the rest of the world, unaffected MENA countries, or matched countries identified through coarsened exact matching. For more detailed information on the matching procedure, refer to the Supplementary Text in the supplementary materials, Table S6 for the list of countries and Fig. S14 for checking the distribution of covariates for target and control groups.
@@
where the parameter $\beta^{p}_{s}$ captures the difference in trends between treated and control groups before the treatment period. The parameter $\lambda$ reflects the overall time trend common to all countries, while $\theta$ measures the effect of the time-varying covariates $X_{it}$. The term $\gamma_{i}$ accounts for country-specific fixed effects, capturing characteristics that are constant over time but vary across countries. The test’s null hypothesis ($\beta^{p}_{s}=0$) indicates no difference in pre-treatment trends between treated and control countries for subgroup $s\in S$. The alternative hypothesis ($\beta^{p}_{s}\neq 0$) implies a violation of the parallel trends assumption. We conduct the tests at a significance level of $\alpha=0.05$. When the null hypothesis is rejected ($p<\alpha$), it suggests that treated and control countries exhibited different trends before treatment, which undermines the validity of the DiD framework. When the alternate hypothesis is rejected ($p\geq\alpha$), it does not prove that the assumption of parallel trends holds, but it provides statistical support for it. We find that the null hypothesis is only rejected for one coefficient in Table 1 (indicated with a dagger) and in 8 coefficients in Table S5 in Supplementary Materials. We report all $p$-values for these tests in Table S7 and Table S8 in the Supplementary Materials.
@@ Acknowledgments
Topic analysis of Arab Spring–related research
To examine the evolution of Arab Spring-related research topics within each country group, we selected articles published after 2011 that mentioned at least one country from the respective group. The resulting corpora include 856, 560, and 356 documents for the GO, CW, and GC groups, respectively, which provide sufficient data for topic modeling. Notably, 201 documents mentioned both GO and CW countries, 148 mentioned both GO and GC countries, and 107 mentioned both CW and GC countries. We applied BERTopic using a count vectorizer (n-grams of length 1–3) with English stop-words. Topics were modeled separately within each group to capture group-specific dynamics, and an additional model was fitted across all documents to identify shared themes. The topics Turkey’s Regional Role, Gulf Monarchies, Refugees, Social Media, Islamist Parties, Women’s Rights, and Oil & Financial Markets, EU policy & migration appear consistently across both approaches, refer to tables S3 and S4. Therefore, in our analysis, we focus on the evolution of these shared topics after the Arab Spring within the three country groups (GO, CW, and GC).
Coarsened Exact Matching for Difference in Difference
To assess the causal impact of the Arab Spring on scholarly attention and research funding, we employed a Difference-in-Differences (DiD) design. A central challenge in such designs is ensuring that treatment (in this case, the target countries) and control groups are sufficiently comparable in the pre-treatment period.
Therefore, we employed coarsened exact matching (CEM) (?) to identify a more suitable control group. We utilized the Python implementation of the CEM (?) at the country level to construct a balanced control group. First, we restricted the data to the pre-event period (years <2011<2011) and computed the mean values of scholarly attention (log-transformed), GDP per capita (log), population (log), and researcher population (log) when needed. Each continuous covariate was standardized into a z-score and then coarsened into five bins using fixed cutpoints (,3,1.5,0.75, 0.75, 1.5, 3,)(-$\infty$$0.75,\,1.5,\,3,\,\infty$). Strata were defined as the joint combination of the bins across all covariates. We retained only those strata containing at least one treated and one control country. The resulting matched set is represented in Table S6.
Supplementary Tables
@@
Government overthrow Count Civil wars Count Governmental changes Count Arab Spring 304 Syria/Libya/Yemen 313 Morocco 127 Egypt/Tunisia 253 Arab Spring 132 Arab Spring 103 Social Media 146 Social Media 50 Social Media 48 Oil & Financial Markets 50 Turkey’s Regional Role 41 Gulf Monarchies 29 Women’s Rights 41 Refugees 14 Palestinian Refugees in Jordan 24 EU Policy & Migration 27 EU Security Agenda & Migration 10 Oil & Financial Markets 14 Turkey’s Regional Role 20 Kuwait 11 Inequality, Employment 15
Table S3: Topics and their frequencies across the three country groups (GO, CW, GC), based on one topic model per group.
Topic Government overthrow Civil wars Governmental changes Arab Spring 256 163 113 Egypt 171 6 4 Tunisia 38 7 7 Libya 13 85 1 Syria 3 66 4 Yemen 1 21 Morocco 13 40 Jordan 3 6 21 Kuwait 4 3 34 Gulf monarchies: Saudi 7 13 23 Russia & Iran 3 17 2 Refugees 5 18 9 Social Media 143 60 36 Oil & Financial Markets 44 15 18 Women’s Rights 43 3 9 Islamist parties 36 4 5 Tourism 11 3 8 Military 10 11 6 Employment, Education, and Economy 10 1 2 EU Policy & Migration 22 8 11 Water Scarcity, Climate Change, and Resources 3 9 3 Turkey’s Regional Role 17 41
Table S4: Frequencies of topics across the three country groups using one unified topic model for all Arab Spring-related research.
@@
Group Attention Type Matched Covariates Matched Countries CEM group (1) Total attention Pre-Arab Spring average log total attention, average log GDP per capita, average log population, average log researcher population Angola, Bulgaria, Bolivia, Cameroon, Costa Rica, Cuba, Dominican Republic, Algeria, Ecuador, Honduras, Haiti, Iraq, Kazakhstan, Kyrgyzstan, Laos, Lebanon, Sri Lanka, Malaysia, Peru, Puerto Rico, Paraguay, Romania, Rwanda, Senegal, El Salvador, Slovenia, Togo, Trinidad and Tobago, Ukraine, Venezuela CEM group (2) Domestic attention Pre-Arab Spring average log domestic attention, average log GDP per capita, average log researcher population Angola, Albania, Bulgaria, Bosnia and Herzegovina, Belarus, Bolivia, Botswana, Cameroon, Colombia, Costa Rica, Cuba, Cyprus, Dominican Republic, Algeria, Ecuador, Estonia, Gabon, Guatemala, Honduras, Croatia, Indonesia, Iraq, Jamaica, Kazakhstan, Cambodia, Lebanon, Sri Lanka, Lithuania, Latvia, North Macedonia, Mozambique, Nicaragua, Peru, Philippines, Puerto Rico, Paraguay, Palestine, Romania, Saudi Arabia, Senegal, El Salvador, Slovakia, Slovenia, Trinidad and Tobago, Ukraine, Uruguay, Venezuela CEM group (3) Foreign attention Pre-Arab Spring average log foreign attention, average log GDP per capita, average log population Angola, Azerbaijan, Bulgaria, Belarus, Bolivia, Cameroon, Colombia, Costa Rica, Cuba, Dominican Republic, Algeria, Ecuador, Guatemala, Honduras, Haiti, Iran, Iraq, Kazakhstan, Kyrgyzstan, Laos, Lebanon, Sri Lanka, Malaysia, Peru, Nicaragua, Philippines, Poland, Puerto Rico, Paraguay, Romania, Senegal, Sierra Leone, El Salvador, Slovenia, Chad, Tajikistan, Trinidad and Tobago, Ukraine, Uzbekistan, Venezuela
Table S6: Coarsened Exact Matching (CEM) country matches across attention types. Shared countries across all three groups are shown in bold.
@@
CEM group(3) 0.2383 0.001 0.007 0.465
Table S7: $p$-values for pre-treatment parallel trend for scholarly attention Models control for log GDP, population, and research population where applicable. Each colored column represents $p$-values from separate regressions comparing the target countries with two control groups: the rest of the world and MENA countries less affected by the Arab Spring.
@@
Table S8: $p$-values for pre-treatment parallel trend test for funding Models control for log GDP, population where applicable. Each colored column represents $p$-values from separate regressions comparing the target countries with three control groups: the rest of the world and MENA countries less affected by the Arab Spring, and the countries matched through the coarsened exact matching
@@
Figure S9: Topic dynamics across Arab Spring country groups. Panels (a–c) show the annual topic counts and panels (d–f) the topic shares for scholarly attention covering different thematic areas between 2011 and 2019. Countries are grouped according to their outcomes following the Arab Spring: government overthrow (a, d), civil war (b, e), and governmental changes (c, f). Topic counts indicate the raw number of publications per year mentioning each theme. Prominent topics include social media, oil and financial markets, EU policy and migration, Islamist parties, Turkey’s regional role, and women’s rights.
@@
Figure S13: Change in scholarly attention to target countries as a function of the pre–Arab Spring yearly average of researcher migration from those countries. A strong positive association is observed (Spearman’s $\rho=0.745$, $p<0.0001$; Pearson’s $\rho=0.781$, $p<0.0001$). Countries such as the USA, Saudi Arabia, and France stand out with both high pre–Arab Spring migration levels and large increases in scholarly attention, while others (e.g., Australia, Iran, Egypt, and India) show disproportionately high attention increases relative to their smaller migration baselines. Most other countries cluster near the origin, reflecting limited migration flows and modest shifts in attention.
Figure S14: Distribution of covariates across treated and control units under coarsened exact matching (CEM). Each column corresponds to one CEM control group specification: (a) total attention, (b) domestic attention, and (c) foreign attention. Rows display the distributions of pre-Arab Spring average log-transformed covariates: attention, GDP, population, and researcher population. Gray bars indicate the matched control units, while red bars show the treated (target) units. The overlapping shapes suggest that the CEM procedure achieves a reasonable balance across key covariates, though some discrepancies remain. Empty panels are marked as “Not applicable,” reflecting variables not included for CEM matching.